an economist <b><i>believes</i></b> there is a linear relationship between the market price of a particular commodity and the number of units <b><i>suppliers</i></b> of the commodity are willing to bring to the marketplace. two sample observations indicate that when the price equals $15 per unit, the weekly supply equals 30000 units, and when the price equals $20 per units, the weekly supply equals 48000 units.

1) Predict the weekly supply if the market price equal $25 per unit.


1
Expert's answer
2021-07-18T08:46:03-0400

1) The supply function is

"\frac{P - 15} {20 - 15} = \frac{Q - 30,000} {48,000 - 30,000},"

3,600×(P - 15) = Q - 30,000,

Qs = 3,600P - 24,000.

The weekly supply if the market price equal $25 per unit is

Qs = 3,600×25 - 24,000 = 66,000.